# Bayesian Intrinsic Groupwise Image Registration: Unsupervised Disentanglement of Anatomy and Geometry

### v1

@@ Abstract
This article presents a general Bayesian learning framework for multi-modal groupwise registration on medical images. The method builds on probabilistic modelling of the image generative process, where the underlying common anatomy and geometric variations of the observed images are explicitly disentangled as latent variables. Thus, groupwise registration is achieved through the solution to Bayesian inference. We propose a novel hierarchical variational auto-encoding architecture to realize the inference procedure of the latent variables, where the registration parameters can be calculated in a mathematically interpretable fashion. Remarkably, this new paradigm can learn groupwise registration in an unsupervised closed-loop self-reconstruction process, sparing the burden of designing complex intensity-based similarity measures. The computationally efficient disentangled architecture is also inherently scalable and flexible, allowing for groupwise registration on large-scale image groups with variable sizes. Furthermore, the inferred structural representations from disentanglement learning are capable of capturing the latent anatomy of the observations with visual semantics. Extensive experiments were conducted to validate the proposed framework, including four datasets from cardiac, brain and abdominal medical images. The results have demonstrated the superiority of our method over conventional similarity-based approaches in terms of accuracy, efficiency, scalability and interpretability.
@@ 1 Introduction
Groupwise image registration aims to find the hidden spatial correspondence that aligns multiple observations. For medical images, when the observations reflect some common anatomy, their intrinsic structural correspondence, which can be independent of the multi-modal imaging acquisition protocol, is of particular interest. However, conventional methods on multi-modal groupwise registration usually rely on intensity-based similarity measures to iteratively optimize the spatial transformations. For instance, multivariate joint entropy or mutual information were proposed as groupwise similarity measures in [1, 2, 3, 4], followed by the more computationally favourable template-based approaches [5, 6]. Nevertheless, devising proper similarity measures and choosing the correct registration hyperparameters for heterogeneous medical images can be tedious and challenging. The heavy computational burden and the instability in registration accuracy may also prevent real-world applications of the conventional groupwise registration algorithms.
@@
In this work, we seek to establish a new unified and interpretable learning framework for large-scale unsupervised multi-modal groupwise registration. Inspired by the recent progress in disentangled representation learning [7, 8], we propose to disentangle the underlying common anatomy and geometric variations from the observed images, which can be regarded as reverting the data generating process of the observations. To this end, a probabilistic generative model is constructed, where the common anatomy and spatial transformations are disentangled as latent variables. Thus, the problem of groupwise registration (recovery of the intrinsic structural correspondence) devolves to the estimation of the latent posterior distribution, which is then solved via variational inference. Besides, since unsupervised disentangled representation learning is impossible without proper inductive biases [9, 10], we have designed a novel hierarchical variational auto-encoding architecture to realize the inference procedure of the latent variables, where the decoder network complies with the equivariance assumption of the imaging process. Accordingly, groupwise registration is learned in an unsupervised closed-loop self-reconstruction process: i) the encoder extracts the intrinsic structural representations from the multi-modal observations, based on which the common anatomy and the spatial correspondence are estimated; ii) the decoder emulates the equivariant image generative process from the common anatomy, and reconstructs the observations via the inverse spatial transformations.
Our contributions can be summarized as follows:
@@
  • We propose a new learning paradigm for multi-modal groupwise registration based on Bayesian inference and disentangled representation learning. Remarkably, we can achieve registration in an unsupervised closed-loop self-reconstruction process, which spares the burden of designing complex image similarity measures.
@@
  • We validated the proposed framework on four publicly available medical image datasets, demonstrating its superiority over similarity-based methods in terms of accuracy, efficiency, scalability and interpretability.
The remainder of the article is organized as follows. Section 2 discusses the related work to this study. Section 3 elucidates the proposed Bayesian inference framework for groupwise registration. Section 4 describes the proposed hierarchical variational auto-encoder architecture for learning and estimation of the latent variables. Section 5 presents the experimental setups and evaluation results of our method on four different public datasets. Section 6 notes implications of the proposed framework and concludes the study. For conciseness, we specify in Table I the major mathematical symbols used in the rest of this paper.
TABLE I: Definition of the major mathematical symbols used in this paper.
Symbol
Description
Symbol
Description
$L$
the number of levels of latent variables
$N$
the number of images in the group
$\bm{Z}$
random variable of the common anatomy
$\bm{U}$
random vector of the image group $\{U_{j}\}_{j=1}^{N}$
$\Omega$
the coordinate space of the common anatomy
$\Omega_{j}$
the $j$-th image coordinate space
$\bm{\phi}$
the set of diffeomorphic maps $\bm{\phi}\triangleq\{\phi_{j}\}_{j=1}^{N}$
$\phi_{j}$
the spatial transformation from $\Omega$ to $\Omega_{j}$
$\bm{z}$
the latent common structural representation
$\bm{u}$
an observed sample of the image group
$\bm{z}^{l}$
the latent common structural representation at the $l$-th level
$\bm{v}^{l}$
the latent stationary velocity fields registering the image group at the $l$-th level
$q^{*}(\bm{z}^{l}\,|\,\bm{u},\bm{v})$
the geometric mean categorical variational distribution of $\bm{z}^{l}$ given $\bm{u}\circ\bm{\phi}\triangleq\{u_{j}\circ\phi_{j}\}_{j=1}^{N}$
$\bm{v}_{j}^{l}$
the stationary velocity field of the $j$-th diffeomorphism at the $l$-th level, $\bm{v}^{l}=\{\bm{v}_{j}^{l}\}_{j=1}^{N}$
$q_{j}^{\diamond}(\bm{z}^{l}\,|\,u_{j},\bm{v}_{j})$
the $j$-th single-view categorical variational distribution of $\bm{z}^{l}$ given $u_{j}\circ\phi_{j}$
$\bm{\mu}_{\bm{v},j}^{l},\bm{\Sigma}_{\bm{v},j}^{l}$
parameters of the Gaussian variational distribution $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l})$
$\widetilde{q}_{j}(\bm{z}^{l}\,|\,u_{j})$
the $j$-th single-view categorical variational distribution of $\bm{z}^{l}$ given $u_{j}$
$\bm{v}_{j}^{+}$
the total velocity field aggregating all $\{\bm{v}_{j}^{l}\}_{l=1}^{L}$ upsampled to the top level
$\bm{\psi}$
parameters of the variational model $\{\bm{\psi}_{j}\}_{j=1}^{N}$
$\bm{\theta}$
parameters of the generative model
@@ 2.1 Groupwise Image Registration
where the transformation $\phi_{j}^{-1}:\Omega_{j}\rightarrow\Omega$ maps spatial locations in the image domain to a common coordinate space $\Omega\subset\mathds{R}^{d}$, and $\mathcal{G}$ is often taken as the group of diffeomorphisms. This concept has motivated the development of population averaging methods that estimate the deformations $\mathcal{G}$ and the template $U_{0}$ simultaneously, based on the observations $\bm{U}$ [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The resultant transformations encode the structural variability and the template provides a statistical representative of the images. In particular, a priori guess of the template as one of the observed images is proposed in [16, 17, 19], with the template updated by the average deformation in each iteration. Joshi et al. [18] and Bhatia et al. [20] proposed the usage of intensity mean image as the template, which avoid reference selection by estimating transformations from the common space. Allassonnière et al. [21] extended the setup to a Bayesian framework, with the template modelled as a linear combination of continuous kernel functions. Recently, Dalca et al. [25], Ding and Niethammer [26] proposed learning-based template estimation methods, using spatial transformation networks to predict diffeomorphisms parameterized by stationary velocity fields [27, 28]. Nevertheless, groupwise registration based on the deformable template assumption could not readily accommodate the multi-modal nature of medical images.
Multi-modality can introduce additional complexity, as the observed images are not related to an intensity template directly through spatial transformations. In this case, based on the common anatomy assumption of multi-modal medical images, we can instead assume an anatomical or structural representation $\bm{Z}$ from which each observed image is generated through the composition of an imaging functional $f_{j}$ and a spatial transformation $\phi_{j}^{-1}$, i.e.
@@
Inspired by inter-modality pairwise registration based on the joint intensity distribution (JID) or particularly the mutual information (MI) [33, 34, 35, 36, 37], initial attempts to realize multi-modal groupwise registration focused on generalizing this information-theoretic approach directly to high-dimensional cases [1, 2, 38, 3, 4]. Later, concerning the curse of dimensionality in estimating the high-dimensional JIDs, template-based groupwise registration revives through probabilistic modelling of the image generative process in Eq. 3 [39, 40, 41, 5, 42, 6]. These methods assumed that the JID is modelled by the marginalization of the joint distribution between the latent template and the observed images. Thus, the spatial transformations were determined by maximum likelihood estimators (MLEs) while the template was predicted via maximum a posteriori (MAP). We direct interested readers to our previous work [6] for detailed elucidation of this maximum-likelihood perspective and its connection to information-theoretic metrics.
@@ 2.3 Multi-Modal Representation Learning
Modern probabilistic generative models [56, 57, 58, 59] construct a general framework for representation learning, as the data formation procedure to which they adapt is optimally suited for an unsupervised manner, and they are powerful in striking a balance between fitting and generalizability by taking into account the uncertainty in characterizing the data and latent distributions.
One family of methods within this scope is variational autoencoders (VAEs) [56], where the paradigm is to approximate the intractable latent distribution by a learnable variational posterior during optimizing an evidence lower bound (ELBO) for the log likelihood of observed data. It provides an information-theoretic way to establish the foundational underpinnings of data encoding for downstream tasks [60]. The multi-modal variants of VAEs additionally imbue the capability of retrieving modality-invariant representation by factorizing it into single-view posteriors (a.k.a. experts) inferred from each modality. A typical factorization specifies the combination of the experts explicitly (e.g. mixtures [61], products [62], mixtures of products [63]) or implicitly (e.g. cross-modal generation [64]), compelling the summarization of the common information from any individual modality.
These approaches, nonetheless, all necessitate originally aligned inputs, such as the image and sentence that describe the same object [61], or multi-modal images that are well-registered [65, 66], which limits their applicability to distorted data. Besides, they often adopt a reductionist stance by assuming an overly simplistic prior. (e.g., the standard Gaussians or Laplacians), failing to respect the structure of latent manifold induced by real-world high-dimensional images, which is crucial for mitigating off-track regularization and vanishing latent dimensions [67, 68], as well as for ensuring suitable decomposition of latent encodings [69]. The mixture-based factorizations of the posterior also suffer from inevitable gap between the ELBO and true likelihood, leading to undesirable suboptimal results [70].
In contrast, our framework embraces advancements in tackling all the limitations above: it allows to learn from distorted images in a unified manner, through explicitly endowing the representations of spatial transformations and the common anatomy with a disentangled nature.
@@ 3 Bayesian Groupwise Registration
Figure 1: The proposed hierarchical framework for Bayesian groupwise registration (3-layer example). Random variables are in circles, deterministic variables are in double circles, and observed variables are shaded. Diamonds denotes network feature maps, and squares represent the variational distributions. (a) Probabilistic graphical model of the generative process. (b) Inference steps #1 that predict the hierarchical velocity fields, where we denote $\widetilde{\bm{q}}_{\bm{z}^{l}}\triangleq\{\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})\}_{j=1}^{N}$ and $\bm{q}_{\bm{v}^{l}}\triangleq\{q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\}_{j=1}^{N}$. (c) Inference steps #2 that predict the common structural representations base on the warped images, where we denote $\bm{q}_{\bm{z}^{l}}^{\diamond}\triangleq\{q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\}_{j=1}^{N}$ and $\bm{q}_{\bm{z}^{l}}^{*}\triangleq q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$. (d) Generation steps that reconstruct the original images. Note that the inference and generation steps form a closed-loop self-reconstruction process.
We propose to formulate the estimation of groupwise registration as a problem of Bayesian inference. Specifically, let $\bm{u}=(u_{j})_{j=1}^{N}$ be a sample of the random vector $\bm{U}$ of the image group. From Eq. 3, each image $u_{j}$ is generated from the latent variable $\bm{z}$ representing the corresponding common anatomy, and the diffeomorphic transformation $\phi_{j}^{-1}$. To ensure its invertibility, we assume that the transformations $\bm{\phi}\triangleq(\phi_{j})_{j=1}^{N}$ is parameterized by stationary velocity fields $\bm{v}=(\bm{v}_{j})_{j=1}^{N}$ [27, 28] such that $\phi_{j}=\exp(\bm{v}_{j})$, and

### v2

# Bayesian Unsupervised Disentanglement of Anatomy and Geometry for Deep Groupwise Image Registration

@@ Abstract
This article presents a general Bayesian learning framework for multi-modal groupwise image registration. The method builds on probabilistic modelling of the image generative process, where the underlying common anatomy and geometric variations of the observed images are explicitly disentangled as latent variables. Therefore, groupwise image registration is achieved via hierarchical Bayesian inference. We propose a novel hierarchical variational auto-encoding architecture to realise the inference procedure of the latent variables, where the registration parameters can be explicitly estimated in a mathematically interpretable fashion. Remarkably, this new paradigm learns groupwise image registration in an unsupervised closed-loop self-reconstruction process, sparing the burden of designing complex image-based similarity measures. The computationally efficient disentangled network architecture is also inherently scalable and flexible, allowing for groupwise registration on large-scale image groups with variable sizes. Furthermore, the inferred structural representations from multi-modal images via disentanglement learning are capable of capturing the latent anatomy of the observations with visual semantics. Extensive experiments were conducted to validate the proposed framework, including four different datasets from cardiac, brain, and abdominal medical images. The results have demonstrated the superiority of our method over conventional similarity-based approaches in terms of accuracy, efficiency, scalability, and interpretability.
@@ 1 Introduction
Groupwise image registration aims to find the hidden spatial correspondence that aligns multiple observations. For medical images, when the observations reflect some common anatomy, their intrinsic structural correspondence, which can be independent of the multi-modal imaging acquisition protocol, is of particular interest. However, conventional methods on multi-modal groupwise registration usually rely on intensity-based similarity measures to iteratively optimise the spatial transformations. For instance, multivariate joint entropy and mutual information were proposed as groupwise similarity measures in [1, 2, 3, 4], followed by more computationally favourable template-based approaches [5, 6]. Nevertheless, devising proper similarity measures and choosing the correct registration hyperparameters for heterogeneous medical images can be tedious and challenging. The high computational burden and the instability in registration accuracy may also prevent real-world applications of conventional groupwise registration algorithms.
@@
In this work, we seek to establish a new unified and interpretable learning framework for large-scale unsupervised multi-modal groupwise registration. Inspired by recent progress in disentangled representation learning [7, 8], we propose to disentangle the underlying common anatomy and geometric variations from the observed images, which can be regarded as reverting the data generating process of the observations. To this end, a probabilistic generative model is constructed, where the common anatomy and spatial transformations are disentangled as latent variables. Thus, the problem of groupwise registration (recovery of the intrinsic structural correspondence) devolves to the estimation of the latent posterior distribution, which is then solved via variational inference. Besides, since unsupervised learning of disentangled representation is impossible without proper inductive biases [9, 10], we have designed a novel hierarchical variational auto-encoding architecture to realise the inference procedure of latent variables, where the decoder network complies with the equivariance assumption of the imaging process. Accordingly, groupwise registration is learnt in an unsupervised closed-loop self-reconstruction process: i) the encoder extracts the intrinsic structural representations from the multi-modal observations, based on which the common anatomy and the spatial correspondence are estimated; ii) the decoder emulates the equivariant image generative process from the common anatomy, and reconstructs the observations via the inverse spatial transformations.
To improve the efficiency and scalability of groupwise registration, particularly in achieving deep groupwise registration with variable (and potentially very large) group sizes for the first time, we make the following contributions:
@@
  • We propose a new learning paradigm for multi-modal groupwise registration based on Bayesian inference and disentangled representation learning. Remarkably, we can achieve registration in an unsupervised closed-loop self-reconstruction process, which spares the burden of designing complex image-based similarity measures.
@@
  • We validated the proposed framework on four publicly available medical image datasets, demonstrating its superiority over similarity-based methods in terms of accuracy, efficiency, scalability, and interpretability.
The remainder of the article is organised as follows. Section 2 discusses the related work to this study. Section 3 elucidates the proposed Bayesian inference framework for groupwise registration. Section 4 describes the proposed hierarchical variational auto-encoder architecture for learning and estimation of the latent variables. Section 5 presents the experimental setups and evaluation results of our method on four different public datasets. Section 6 notes implications of the proposed framework and concludes the study. For conciseness, in Table I we specify the main mathematical symbols used in the rest of this paper.
TABLE I: Definition of the main mathematical symbols used in this paper.
Symbol Description Symbol Description
$L$ the number of levels of latent variables $N$ the number of images in the group
$\bm{Z}$ random variable of the common anatomy $\bm{U}$ random vector of the image group $\{U_{j}\}_{j=1}^{N}$
$\Omega$ the coordinate space of the common anatomy $\Omega_{j}$ the $j$-th image coordinate space
$\bm{\phi}$ the set of diffeomorphic maps $\bm{\phi}\triangleq\{\phi_{j}\}_{j=1}^{N}$ $\phi_{j}$ the spatial transformation from $\Omega$ to $\Omega_{j}$
$\bm{z}$ the latent common structural representation $\bm{u}$ an observed sample of the image group
$\bm{z}^{l}$ the latent common structural representation at the $l$-th level $\bm{v}^{l}$ the latent stationary velocity fields registering the image group at the $l$-th level
$q^{*}(\bm{z}^{l}\,|\,\bm{u},\bm{v})$ the geometric mean variational distribution of $\bm{z}^{l}$ given $\bm{u}\circ\bm{\phi}\triangleq\{u_{j}\circ\phi_{j}\}_{j=1}^{N}$ $\bm{v}_{j}^{l}$ the stationary velocity field of the $j$-th diffeomorphism at the $l$-th level, $\bm{v}^{l}=\{\bm{v}_{j}^{l}\}_{j=1}^{N}$
$q_{j}^{\diamond}(\bm{z}^{l}\,|\,u_{j},\bm{v}_{j})$ the $j$-th single-view variational distribution of $\bm{z}^{l}$ given $u_{j}\circ\phi_{j}$ $\bm{\mu}_{\bm{v},j}^{l},\bm{\Sigma}_{\bm{v},j}^{l}$ parameters of the Gaussian variational distribution $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l})$
$\widetilde{q}_{j}(\bm{z}^{l}\,|\,u_{j})$ the $j$-th single-view variational distribution of $\bm{z}^{l}$ given $u_{j}$ $\bm{v}_{j}^{+}$ the total velocity field aggregating all $\{\bm{v}_{j}^{l}\}_{l=1}^{L}$ upsampled to the top level
$\bm{\psi}$ parameters of the variational model $\{\bm{\psi}_{j}\}_{j=1}^{N}$ $\bm{\theta}$ parameters of the generative model
@@ 2.1 Groupwise Image Registration
where the transformation $\phi_{j}^{-1}:\Omega_{j}\rightarrow\Omega$ maps spatial locations in the image domain to a common coordinate space $\Omega\subset\mathds{R}^{d}$, and $\mathcal{G}$ is often taken as the group of diffeomorphisms. This concept has motivated the development of population averaging methods that estimate the deformations $\mathcal{G}$ and the template $U_{0}$ simultaneously, based on the observations $\bm{U}$ [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The resultant transformations encode the structural variability and the template provides a statistical representative of the images. In particular, a priori guess of the template as one of the observed images was proposed in [16, 17, 19], with the template updated by the average deformation in each iteration. Joshi et al. [18] and Bhatia et al. [20] proposed using the intensity mean image as a template, avoiding reference selection by estimating transformations from the common space. Allassonnière et al. [21] extended the setup to a Bayesian framework, with the template modelled as a linear combination of continuous kernel functions. Recently, Dalca et al. [25], Ding and Niethammer [26] proposed learning-based template estimation methods, using spatial transformation networks to predict diffeomorphisms parameterized by stationary velocity fields [27, 28]. Nevertheless, groupwise registration based on the deformable template assumption could not readily accommodate the multi-modal nature of medical images.
Multi-modality can introduce additional complexity, as the observed images are not related to an intensity template directly through spatial transformations. In this case, based on the common anatomy assumption of multi-modal medical images, we can instead assume an anatomical or structural representation $\bm{Z}$ from which each observed image is generated through the composition of an imaging functional $f_{j}$ and a spatial transformation $\phi_{j}^{-1}$, i.e.,
@@
Inspired by inter-modality pairwise registration based on the joint intensity distribution (JID) or particularly the mutual information (MI) [33, 34, 35, 36, 37], initial attempts to realise multi-modal groupwise registration focused on generalising this information-theoretic approach directly to high-dimensional cases [1, 2, 38, 3, 4]. Later, concerning the curse of dimensionality in estimating high-dimensional JIDs, template-based groupwise registration revives through probabilistic modelling of the image generative process in Eq. 3 [39, 40, 41, 5, 42, 6]. These methods assumed that the JID is modelled by the marginalisation of the joint distribution between the latent template and the observed images. Thus, the spatial transformations were determined by maximum likelihood estimators (MLEs) while the template was predicted via maximum a posteriori (MAP). We direct interested readers to our previous work [6] for detailed elucidation of this maximum-likelihood perspective and its connection to information-theoretic metrics.
@@ 2.3 Multi-Modal Representation Learning
Modern probabilistic generative models [58, 59, 60] construct a general framework for representation learning, since the data formation procedure to which they adapt is optimally suited for an unsupervised manner, and are powerful in striking a balance between fitting and generalisability taking into account the uncertainty in characterising the data and latent distributions.
One family of methods within this scope is variational autoencoders (VAEs) [58], where the paradigm is to approximate the intractable latent distribution by a learnable variational posterior during the optimisation of an evidence lower bound (ELBO) of the log likelihood of observed data. It provides an information-theoretic way to establish the foundational underpinnings of data encoding for downstream tasks [61]. The multi-modal variants of VAEs additionally imbue the capability of retrieving modality-invariant representation by factorising it into single-view posteriors (a.k.a. experts) inferred from each modality. A typical factorisation specifies the combination of the experts explicitly (e.g. mixtures [62], products [63], mixtures of products [64]) or implicitly (e.g. cross-modal generation [65]), compelling the summarisation of the common information from any individual modality.
These approaches, nonetheless, all necessitate originally aligned inputs, such as the image and sentence that describe the same object [62], or multi-modal images that are well registered [66, 67], which limits their applicability to distorted data. Besides, they often adopt a reductionist stance by assuming an overly simplistic prior. (e.g., the standard Gaussians or Laplacians), failing to respect the structure of latent manifold induced by real-world high-dimensional images, which is crucial for mitigating off-track regularisation and vanishing latent dimensions [68, 69], as well as for ensuring suitable decomposition of latent encodings [70]. The mixture-based factorisation of the posterior also suffers from inevitable gap between the ELBO and the true likelihood, leading to undesirable suboptimal results [71].
In contrast, our framework embraces advancements in tackling all the limitations above: It allows learning from distorted images in a unified manner, through explicitly endowing the representations of spatial transformations and the common anatomy with a disentangled nature.
@@ 3 Bayesian Groupwise Registration
Figure 1: The proposed hierarchical framework for Bayesian groupwise registration (3-layer example). Random variables are in circles, deterministic variables are in double circles, and observed variables are shaded. Diamonds denote network feature maps, and squares represent variational distributions. (a) Probabilistic graphical model of the generative process. (b) Inference steps #1 that predict the hierarchical velocity fields, where we denote $\widetilde{\bm{q}}_{\bm{z}^{l}}\triangleq\{\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})\}_{j=1}^{N}$ and $\bm{q}_{\bm{v}^{l}}\triangleq\{q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})\}_{j=1}^{N}$. (c) Inference steps #2 that predict the common structural representations base on the warped images, where we denote $\bm{q}_{\bm{z}^{l}}^{\diamond}\triangleq\{q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})\}_{j=1}^{N}$ and $\bm{q}_{\bm{z}^{l}}^{*}\triangleq q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$. (d) Generation steps that reconstruct the original images. Note that the inference and generation steps form a closed-loop self-reconstruction process.
The fundamental principle driving the proposed approach is to disentangle the underlying common anatomy and geometric variations as latent representations w.r.t. the decomposition of two symmetry groups acting independently in the observational image space. More details on the underlying connection between disentangled representations and symmetry groups can be found in Section E.1 of the Supplementary Material. To this end, we propose to formulate the estimation of these latent variables as a problem of Bayesian inference. Specifically, let $\bm{u}=(u_{j})_{j=1}^{N}$ be a sample of the random vector $\bm{U}$ of the image group. From Eq. 3, each image $u_{j}$ is generated from the latent variable $\bm{z}$ representing the corresponding common anatomy, and the diffeomorphic transformation $\phi_{j}^{-1}$. To ensure its invertibility, we assume that the transformations $\bm{\phi}\triangleq(\phi_{j})_{j=1}^{N}$ is parameterized by stationary velocity fields $\bm{v}=(\bm{v}_{j})_{j=1}^{N}$ [27, 28] such that $\phi_{j}=\exp(\bm{v}_{j})$, and